# R-PINN: Recovery-type a-posteriori estimator enhanced adaptive PINN

### v1

@@ 4 Numerical experiments
In this section, we present numerical experiments demonstrating the effectiveness of our proposed algorithm. Our evaluation compares three different kinds of PINNs: the Recovery-type estimator enhanced PINN (R-PINN), Failure-informed PINN (FI-PINN) [20], and standard PINN as a baseline.
@@ 4.1 Experimental setup
To ensure a fair comparison, we maintain identical parameters and training processes across all methods for the same number of iterations, varying only the number and distribution of adaptive collocation points. The recovery-type a-posteriori estimation is implemented using FEALPy [37].
@@
All experiments employ a fully-connected neural network with $7$ hidden layers, each containing 20 neurons with tanh for the activation function unless otherwise specified. The training process comprises two phases: an initial pre-training phase utilizing the L-BFGS optimizer with a learning rate of $0.1$, conducted for a problem-dependent number of epochs, followed by an adaptive training phase consisting of $M$ adaptive iterations. Each iteration is constrained to $5000$ epochs while employing the same optimization method.
@@
For collocation points, we use $N_{1}=2000$ background collocation points and $N_{2}=100$ adaptive collocation points for R-PINN unless specified, while FI-PINN requires $N_{2}=500$ adaptive collocation points to achieve convergence unless specified. For the standard PINN, only $N_{1}=2000$ background collocation points are used during training, without incorporating additional adaptive collocation points. However, the training process remains the same as that of the two adaptive PINNs, with each adaptive iteration also consisting of 5000 epochs. All methods utilize $200$ boundary collocation points. Notably, our approach achieves comparable or superior accuracy despite using fewer total epochs and adaptive collocation points than related studies. The larger $N_{2}$ value for FI-PINN is necessary as it fails to achieve adaptivity with fewer points. Furthermore, while our method maintains a fixed number of adaptive collocation points, FI-PINN accumulates points throughout the adaptive iteration process. To evaluate the accuracy of our method, we employ two error metrics: the relative $L^{2}$ error,
$\frac{\lVert\hat{u}{\tilde{\Theta}}-u\rVert_{2}}{\lVert u\rVert_{2}},$ (4.1)
@@
and the $L^{\infty}$ error, defined as $\lVert\hat{u}{\tilde{\Theta}}-u\rVert_{\infty}$. Here, $\hat{u}_{\tilde{\Theta}}$ denotes the numerical solution obtained by our method, and $u$ represents the exact solution.
4.2 Poisson equation
Consider the following Poisson equation:
$\begin{split}-\Delta u(x,y)=&f(x,y),\quad(x,y)\ \text{in}\ \Omega,\\ u(x,y)=&g(x,y),\quad(x,y)\ \text{on}\ \partial\Omega,\end{split}$ (4.2)
where the domain $\Omega$ is $=[0,1]\times[0,1]$, and the true solution is
$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}.$ (4.3)
@@ 4.2 Poisson equation
The solution exhibits a sharp peak near the point $(0.5,0.5)$, which is challenging for standard PINN to capture accurately. In this experiment, we pre-train the network for $10000$ iterations and set the maximum number of adaptive iterations to $M=5$.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $3.386$ $0.0990$ $100$ $2.717$ $0.6541$ $500$ $0.809$ $0.6270$
$2$ $0.732$ $0.0282$ $100$ $1.514$ $0.4593$ $1000$ $0.689$ $0.6180$
$3$ $0.125$ $0.0260$ $100$ $1.157$ $0.3095$ $1500$ $0.650$ $0.5683$
$4$ $0.285$ $0.0088$ $100$ $0.294$ $0.0686$ $2000$ $0.879$ $0.5689$
$5$ $0.139$ $0.0086$ $100$ $0.068$ $0.0335$ $2500$ $0.682$ $0.5570$
@@
Table 1: The relative $L^{2}$ error and $L^{\infty}$ error by two methods solving Poisson equation.
@@
Fig. 4: The distribution of adaptive collocation points in the first $2$ adaptive iterations with two adaptive PINN methods.
Fig. 4 illustrates the distribution of adaptive collocation points by two adaptive methods in the first two adaptive iterations. It is evident that all adaptive collocation points of our model converge to the neighborhood of the point $(0.5,0.5)$, while adaptive collocation points of FI-PINN are more dispersed. This shows the strong effectiveness of our algorithm in handling problems with significant local variations. Table 1 compares the relative $L^{2}$ error and $L^{\infty}$ error among R-PINN, FI-PINN and standard PINN during the adaptive iterations where $N_{\text{iter}}$ represents the number of adaptive iterations. The relative $L^{2}$ error for R-PINN reaches $0.139$, and the $L^{\infty}$ error reaches $0.0086$. In contrast, the errors obtained by FI-PINN decrease slowly at the beginning, but drop fast later, reaching the relative $L^{2}$ error of $0.068$ and $L^{\infty}$ error of $0.0335$. It can be observed that R-PINN can converge faster than FI-PINN. While, R-PINN can reach a smaller relative $L^{2}$ error in this case with the exponentially accumulated adaptive collocation points, which is one of its advantage. As for the normal PINN, the errors obtained show little to no decrease, indicating that the normal PINN fails to converge effectively. Similar results are observed in Fig. 6, where the lowest relative $L^{2}$ error for our algorithm is below $1\times 10^{-1}$. In the left panel of Fig. 5, we show the absolute errors throughout the entire domain, clearly demonstrating that the errors are much smaller with R-PINN. The middle panel of Fig. 5 illustrates the cross-section at $y=0.5$. The predicted curve using R-PINN fits the true solution very well, while the solution by FI-PINN has a little obvious error at the foot of the peak. The final solutions with three methods are plotted in the right of Fig. 5.
@@
Fig. 5: The prediction err (left), the cross section at $y=0.5$ (middle) and the predicted solution of three methods solving Poisson equation.
Fig. 6: The relative $L^{2}$ errors of three methods in the optimization process of solving Poison equation.
@@ 4.3 Burgers’ equation
$\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\frac{0.01}{\pi}\frac{\partial^{2}u}{\partial x^{2}},\\ &u(x,0)=-\sin(\pi x),\\ &u(t,-1)=u(t,1)=0\end{split}$ (4.4)
where the domain is $(x,t)\in[-1,1]\times[0,1]$. The solution to this Burgers’ equation exhibits a sharp mutation along $x=0$, making it prone to errors and offsets. For this test, we pre-train the network for $5000$ iterations and use $M=4$ steps of adaptive processes.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $0.0086$ $0.0735$ $100$ $0.2278$ $1.8142$ $200$ $0.0136$ $0.1583$
$2$ $0.0107$ $0.0964$ $100$ $0.0131$ $0.1263$ $400$ $0.0144$ $0.1685$
$3$ $0.0071$ $0.0673$ $100$ $0.0064$ $0.0576$ $600$ $0.0174$ $0.2099$
$4$ $0.0059$ $0.0439$ $100$ $0.0089$ $0.0846$ $800$ $0.0197$ $0.2418$
Table 2: The relative $L^{2}$ error and $L^{\infty}$ error by three methods solving Burgers’ equation.
@@
Fig. 7: The distribution of adaptive collocation points in the first $2$ adaptive iterations with two adaptive PINN methods solving Burgers’ equation.
The adaptive collocation points generated during the first two adaptive training processes by two adaptive PINN method are illustrated in Fig. 7. We observe that the adaptive collocation points successfully cluster around $x=0$ with both methods, where the error is largest (as seen in Fig. 8). This clustering significantly improves the local error. The exact relative $L^{2}$ error and $L^{\infty}$ error are listed in Table 2. As shown, the final relative $L^{2}$ error and $L^{\infty}$ error for R-PINN are $0.0059$ and $0.0439$, respectively, while those for FI-PINN are $0.0089$ and $0.0846$. The prediction err of standard PINN keeps at the save level during the training, where relative $L^{2}$ error and $L^{\infty}$ reach $0.0197$ and $0.2418$ in the final. Similar to the experiment of Poisson equation, it takes a long iterations for FI-PINN before the error jumps quickly, while R-PINN can reduce the errors at once. The errors obtained by R-PINN and FI-PINN are one order of magnitude lower than those obtained by the standard PINN. Such behaviour can also be observed in the Fig. 10, where the relative $L^{2}$ errors are illustrated specifically, showing that the lowest error can reach below $1\times 10^{-2}$. We also show the absolute errors across the entire domain and the cross-section at $t=1$ in Fig. 8. The predicted solutions of two adaptive methods are illustrated in Fig. 9.
@@
Fig. 8: The prediction err (left) and the cross section at $t=1$ (right) of three methods solving Burgers’ equation.
@@
Fig. 10: The relative $L^{2}$ errors of three methods in the optimization process of solving Burgers’ equation.
@@ 4.4 Two-dimensional problem with two peaks
$\begin{split}-\nabla\cdot[u(x,y)\nabla(x^{2}+y^{2})]+\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in\Omega,\\ u(x,y)=b(x,y),\quad(x,y)\in\partial\Omega,\end{split}$ (4.5)
@@
$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}+e^{-1000[(x+0.5)^{2}+(y+0.5)^{2}]},$ (4.6)
which has two peaks at $(-0.5,-0.5)$ and $(0.5,0.5)$. This test mainly can show the ability of R-PINN to deal with multiple peaks. For this experiment, we pre-train our network for the period of $10000$ iterations and set the maximum number of adaptive iterations as $M=4$. Since two peaks problem is challenging for neural network to approximate, we use $5$ hidden layers with $64$ neurons in each layer to enhance its capability. Moreover, we use $N_{2}=150$ adaptive collocation points for R-PINN, and $N_{2}=750$ added adaptive collocation points for FI-PINN.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $0.6778$ $0.1228$ $150$ $0.4372$ $0.4845$ $750$ $0.7912$ $0.6651$
$2$ $0.1202$ $0.0198$ $150$ $0.5220$ $0.3429$ $1500$ $0.8020$ $0.6475$
$3$ $0.1080$ $0.0259$ $150$ $0.4773$ $0.2233$ $2250$ $0.8012$ $0.6483$
$4$ $0.0840$ $0.0127$ $150$ $0.3969$ $0.2263$ $3000$ $0.8342$ $0.6625$
Table 3: The relative $L^{2}$ error and $L^{\infty}$ error by two methods solving two-dimensional problem with two peaks.
@@
Fig. 11: The distribution of adaptive collocation points in the first two adaptive iterations of two adaptive methods solving the two-dimensional problem with two peaks.
@@
Fig. 12: The predicted err (left), the cross section at $y=-0.5$ (middle) and the cross section at $y=0.5$ (right) of three methods solving the two-dimensional problem with two peaks.
Fig. 13: The predicted solution (R-PINN at the left, FI-PINN at the middle and standard PINN at the right) of the two-dimensional problem with two peaks.
Fig. 14: The relative $L^{2}$ errors of three methods in the optimization process of solving the two-dimensional problem with two peaks.
We present the distribution of adaptive collocation points of two adaptive methods across the first two adaptive iterations in Fig. 11. It is evident that these points converge around the vicinities of $(0.5,0.5)$ and $(-0.5,-0.5)$ with R-PINN, demonstrating the capability of our algorithm to effectively handle problems characterized by multiple regions of abrupt changes. However, the distribution of adaptive collocation points with FI-PINN seems not so effective to deal with the two peaks. The reason of it might be a truncated Gaussian distribution is used, so that problems of more than one regions with large error are difficult to handle. This problem is also reflected in the training process. Table 3 compares the relative $L^{2}$ and $L^{\infty}$ errors among R-PINN, FI-PINN and standard PINN during the adaptive iterations. For R-PINN, the relative $L^{2}$ error reduces to $0.0840$, and the $L^{\infty}$ error decreases to $0.0127$. In contrast, the relative $L^{2}$ error associated with FI-PINN can only reach $0.3969$ and the $L^{\infty}$ can get to $0.2263$, which is just a little bit lower than those with standard PINN. Similar trends are observed in Fig. 14, where the lowest relative $L^{2}$ error for R-PINN falls below $1\times 10^{-1}$.
Fig. 12 provides additional insights. The left panel illustrates the absolute errors across the entire domain, revealing significantly smaller errors for R-PINN. The middle and right panels display cross-sectional views at $y=-0.5$ and $y=0.5$, respectively, showing that the predictions from R-PINN align closely with the exact solution, whereas there are large errors near two peaks with FI-PINN and the standard PINN can only approximate the features. Finally, the numerical solutions obtained by the three methods are visualized in Fig. 13.

### v2

@@ 4 Numerical experiments
In this section, we present numerical experiments demonstrating the effectiveness of our proposed algorithm. Our evaluation compares three kinds of PINNs: recovery-type estimator enhanced PINN (R-PINN), failure-informed PINN (FI-PINN) [9] and residual-based attention in PINN (RBA-PINN) [1]. Among these methods, the first two adopt the adaptive sampling strategy, while RBA-PINN employs a strategy of adaptive weighting. For the gradient reconstruction in R-PINN, we employ the weighted averaging approach throughout all experiments. It is observed that the other two reconstruction methods achieve similarly comparable performance.
@@ 4.1 Experimental setup
To ensure a fair comparison, we maintain identical parameters and training processes across all methods for the same number of iterations, varying only the number and distribution of adaptively distributed points. The recovery-type a-posteriori estimation is implemented using FEALPy [34].
@@
All experiments employ a fully-connected neural network with $7$ hidden layers, each containing 20 neurons with tanh for the activation function unless otherwise specified. The training process comprises two phases: an initial pre-training phase utilizing the L-BFGS optimizer with a learning rate of $0.1$, conducted for a problem-dependent number of epochs, followed by an adaptive training phase consisting of $M$ adaptive iterations. Each iteration is constrained to $5000$ epochs while employing the same optimization method unless otherwise specified.
Fig. 3: Illustration of a $10\times 10$ regular diagonal triangulation.
@@
For the choice of parameters, we set $\epsilon=0.02$ for R-PINN. The triangulation of the domain adopted in the method is a uniform diagonal mesh composed of right triangles. An example with a $10\times 10$ mesh is illustrated in Fig. 3. Unless otherwise specified, we use a $50\times 50$ mesh as the default configuration in the following experiments. For collocation points, we use $N_{1}=2000$ background collocation points and $N_{2}=100$ adaptively distributed points for R-PINN unless specified. For the RBA-PINN, only $N_{1}=2000$ background collocation points are used during training, without incorporating additional adaptively distributed points unless specified. However, the training process remains the same as that of the two adaptive sampling PINNs. All methods utilize $200$ boundary collocation points. Notably, our approach achieves comparable or superior accuracy despite using fewer total epochs and adaptively distributed points than related studies. The larger $N_{2}$ value for FI-PINN is necessary as it fails to achieve adaptivity with fewer points. Furthermore, while our method maintains a fixed number of adaptively distributed points, FI-PINN accumulates points throughout the adaptive iteration process. To evaluate the accuracy of our method, we employ two error metrics: the relative $L^{2}$ error,
$\frac{\lVert\hat{u}_{\bm{\Theta}}-u\rVert_{2}}{\lVert u\rVert_{2}},$
@@
and the $L^{\infty}$ error, defined as $\lVert\hat{u}_{\bm{\Theta}}-u\rVert_{\infty}$. Here, $\hat{u}_{\bm{\Theta}}$ denotes the numerical solution obtained by our method, and $u$ represents the exact solution.
4.2 Poisson’s equation
Consider the following Poisson’s equation:
$\begin{split}-\Delta u(x,y)=&f(x,y),\quad(x,y)\ \text{in}\ \Omega,\\ u(x,y)=&g(x,y),\quad(x,y)\ \text{on}\ \partial\Omega,\end{split}$
where the domain $\Omega=[0,1]\times[0,1]$, and the true solution is
$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}.$
@@ 4.2 Poisson’s equation
The solution exhibits a sharp peak near the point $(0.5,0.5)$, which is challenging for vanilla PINN to capture accurately. In this experiment, the network is pre-trained for 15,000 epochs. The number of background collocation points is set to $N_{1}=4000$, and the maximum number of adaptive iterations is specified as $M=4$. Each adaptive iteration consists of 10,000 epochs instead of 5,000, due to the difficulty of approximating this particular example, which features a solution with a sharp peak.
R-PINN FI-PINN RBA-PINN
$N_{\text{iter}}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$
1 0.1695 0.0259 120 0.4166 0.2862 500 0.3042 0.2040
2 0.0673 0.0177 120 0.1656 0.0449 1000 0.3360 0.1880
3 0.0497 0.0103 120 0.1405 0.0160 1500 0.2781 0.1828
4 0.0277 0.0015 120 0.0530 0.0086 2000 0.2741 0.1791
@@
Table 1: The relative $L^{2}$ and $L^{\infty}$ error obtained by three different PINNs solving Poisson’s equation.
@@
Fig. 4: The distribution of adaptively distributed points in the first 22 adaptive iterations with two adaptive PINN methods.
Fig. 4 illustrates the distribution of adaptively distributed points by two adaptive sampling methods in the first two adaptive iterations. It is evident that all adaptively distributed points of our model converge to the neighborhood of the point $(0.5,0.5)$, while adaptively distributed points of FI-PINN are more dispersed. This shows the strong effectiveness of our algorithm in handling problems with significant local variations. Table 1 compares the relative $L^{2}$ error and $L^{\infty}$ error among R-PINN, FI-PINN and RBA-PINN during the adaptive iterations where $N_{\text{iter}}$ represents the number of adaptive iterations. The relative $L^{2}$ error for R-PINN reaches 0.0277, and the $L^{\infty}$ error reaches 0.0015. In contrast, the errors obtained by FI-PINN decrease slowly at the beginning, but drop fast later, reaching the relative $L^{2}$ error of 0.0530 and $L^{\infty}$ error of 0.0086. It can be observed that R-PINN can converge faster than FI-PINN. R-PINN can also reach a smaller relative $L^{2}$ error even in the case where the adaptively distributed points of FI-PINN are continuously accumulated. As for RBA-PINN, it appears to have a difficulty in capturing the local features of the solution accurately. Similar results are observed in Fig. 6, where the lowest relative $L^{2}$ error for our algorithm is below $5\times 10^{-2}$. In the left panel of Fig. 5, we show the absolute errors throughout the entire domain, clearly demonstrating that the errors are much smaller with R-PINN. The middle panel of Fig. 5 illustrates the cross-sections at $y=0.5$. Both R-PINN and FI-PINN are able to accurately approximate the peak-shaped solution. In contrast, although RBA-PINN identifies the local feature, its accuracy is limited by the insufficient number of collocation points in the corresponding region. The final solutions by three different PINNs are plotted in the right of Fig. 5.
@@
Fig. 5: The absolute errors across the domain (left), the cross-sections at $y=0.5$ (middle) and the predicted solutions of three adaptive methods solving Poisson’s equation.
Fig. 6: The relative $L^{2}$ error by three different PINNs in the optimization process of solving Poisson’s equation.
@@ 4.3 Burgers’ equation
$\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\frac{0.001}{\pi}\frac{\partial^{2}u}{\partial x^{2}},\\ &u(x,0)=-\sin(\pi x),\\ &u(t,-1)=u(t,1)=0\end{split}$
where the domain is $(x,t)\in[-1,1]\times[0,1]$. The solution to this Burgers’ equation exhibits an extremely sharp mutation along $x=0$, which makes the problem highly susceptible to numerical errors and local deviations. In this test, we pre-train the network with $N_{1}=50,000$ background collocation points for $15,000$ iterations and use $M=4$ steps of adaptive processes.
R-PINN FI-PINN RBA-PINN
$N_{\text{iter}}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$
1 0.0166 0.3260 500 0.2106 1.8845 600 0.1692 1.9880
2 0.0169 0.4087 500 0.1997 1.9430 1200 0.1692 1.9881
3 0.0122 0.2078 500 0.1926 1.9593 1800 0.1693 1.9881
4 0.0048 0.0920 500 0.1928 1.9471 2400 0.1693 1.9881
Table 2: The relative $L^{2}$ and $L^{\infty}$ error obtained by three different PINNs solving Burgers’ equation.
@@
Fig. 7: The distribution of adaptively distributed points in the first 22 adaptive iterations with two adaptive PINN methods solving Burgers’ equation.
The adaptively distributed points generated during the first two adaptive training processes by two adaptive sampling PINN methods are illustrated in Fig. 7. We observe that the adaptively distributed points successfully cluster around $x=0$ with both methods, where the error is largest (as seen in Fig. 8). This clustering significantly improves the local error. The exact relative $L^{2}$ error and $L^{\infty}$ error are listed in Table 2. As shown, the final relative $L^{2}$ error and $L^{\infty}$ error for R-PINN are 0.0048 and 0.0920, respectively, while those for FI-PINN are 0.1928 and 1.9471. In contrast, the final relative $L^{2}$ and $L^{\infty}$ errors of RBA-PINN are 0.1693 and 1.9881, respectively. Both FI-PINN and RBA-PINN fail to accurately capture the shock, whereas R-PINN exhibits significantly lower errors than the other two methods. Such behaviour can also be observed in Fig. 10, where the relative $L^{2}$ error is illustrated specifically, showing that the lowest error can reach below $1\times 10^{-2}$. We also show the absolute errors across the entire domain and the cross-sections at $t=1$ in Fig. 8. The predicted solutions of two adaptive sampling methods are illustrated in Fig. 9.
@@
Fig. 8: The absolute errors across the domain (left) and the cross-sections at $t=1$ (right) of three different PINNs solving Burgers’ equation.
@@
Fig. 10: The relative $L^{2}$ error by three different PINNs in the optimization process of solving Burgers’ equation.
@@ 4.4 Two-dimensional problem with two peaks
$\begin{split}-\nabla\cdot[u(x,y)\nabla(x^{2}+y^{2})]+\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in\Omega,\\ u(x,y)=b(x,y),\quad(x,y)\in\partial\Omega,\end{split}$
@@
$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}+e^{-1000[(x+0.5)^{2}+(y+0.5)^{2}]},$
which has two peaks at (0.5,0.5)(-0.5,-0.5) and $(0.5,0.5)$. This test mainly can show the ability of R-PINN to deal with multiple peaks. For this experiment, we pre-train our network for the period of $10000$ iterations and set the maximum number of adaptive iterations as $M=4$. Since two peaks problem is challenging for neural network to approximate, we use $5$ hidden layers with $64$ neurons in each layer to enhance its capability. Moreover, we use $N_{2}=150$ adaptively distributed points for R-PINN, and $N_{2}=750$ added adaptively distributed points for FI-PINN.
R-PINN FI-PINN RBA-PINN
$N_{\text{iter}}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$
1 0.6778 0.1228 150 0.4372 0.4845 750 0.4045 0.5299
2 0.1202 0.0198 150 0.5220 0.3429 1500 0.3764 0.4885
3 0.1080 0.0259 150 0.4773 0.2233 2250 0.3578 0.4646
4 0.0840 0.0127 150 0.3969 0.2263 3000 0.3474 0.4507
Table 3: The relative $L^{2}$ and $L^{\infty}$ error obtained by three different PINNs solving the two-dimensional problem with two peaks.
@@
Fig. 11: The distribution of adaptively distributed points in the first two adaptive iterations of two adaptive sampling methods solving the two-dimensional problem with two peaks.
@@
Fig. 12: The absolute errors across the domain (left), the cross-sections at y=0.5y=-0.5 (middle) and the cross-sections at $y=0.5$ (right) of three different PINNs solving the two-dimensional problem with two peaks.
Fig. 13: The predicted solution (R-PINN at the left, FI-PINN at the middle and RBA-PINN at the right) of the two-dimensional problem with two peaks.
Fig. 14: The relative $L^{2}$ error by three different PINNs in the optimization process of solving the two-dimensional problem with two peaks.
We present the distribution of adaptively distributed points of two adaptive sampling methods across the first two adaptive iterations in Fig. 11. It is evident that these points converge around the vicinities of $(0.5,0.5)$ and (0.5,0.5)(-0.5,-0.5) with R-PINN, demonstrating the capability of our algorithm to effectively handle problems characterized by multiple regions of abrupt changes. However, the distribution of adaptively distributed points with FI-PINN seems not so effective to deal with the two peaks. The reason of it might be a truncated Gaussian distribution is used, so that problems of more than one regions with large error are difficult to handle. This problem is also reflected in the training process. Table 3 compares the relative $L^{2}$ and $L^{\infty}$ errors among R-PINN, FI-PINN and RBA-PINN during the adaptive iterations. For R-PINN, the relative $L^{2}$ error reduces to 0.0840, and the $L^{\infty}$ error decreases to 0.0127. In contrast, the relative $L^{2}$ error associated with FI-PINN can only reach 0.3969 and the $L^{\infty}$ error can get to 0.2263, while RBA-PINN achieves a lower relative $L^{2}$ error 0.3474 but a larger $L^{\infty}$ error 0.4507. Similar trends are observed in Fig. 14, where the lowest relative $L^{2}$ error for R-PINN falls below $1\times 10^{-1}$.
Fig. 12 provides additional insights. The left panel illustrates the absolute errors across the entire domain, revealing significantly smaller errors for R-PINN. The middle and right panels display cross-sectional views at y=0.5y=-0.5 and $y=0.5$, respectively, showing that the predictions from R-PINN align closely with the exact solution, whereas there are large errors near two peaks for FI-PINN and RBA-PINN. Finally, the numerical solutions obtained by the three methods are visualized in Fig. 13.
R-PINN FI-PINN RBA-PINN
$N_{\text{iter}}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$ $N_{2}$ Relative $^{2} LL^{\infty}$
1 0.0572 0.0761 100 0.1110 0.1590 500 0.1153 0.0458
2 0.0076 0.0098 100 0.0317 0.0491 1000 0.1140 0.0453
3 0.0076 0.0099 100 0.0317 0.0490 1500 0.1137 0.0452
4 0.0076 0.0099 100 0.0270 0.0418 2000 0.1135 0.0452
Table 4: The relative $L^{2}$ and $L^{\infty}$ error obtained by three different PINNs solving the wave equation.